Can you help me solve this equation?
(a + 4)/(b) = (a - 4)/(b) + (b + 4)/(a)
a = ?

Answer by scott8148(6628)

multiplying by ab gives a(a+4)=a(a-4)+b(b+4)

subtracting a(a-4) gives a(a+4)-a(a-4)=b(b+4) __ a((a+4-(a-4))=b(b+4) __ a(8)=b(b+4)

dividing by 8 gives a=(b^2+4b)/8